If a trolley of mass 5kg moves at 4.m/s east across a frictionless horizontal surface and a box of mass 1,5 kg is dropped onto the trolley then the trolley and the box continues to move in the same direction.calculate the impulse of the trolley.



1
Expert's answer
2022-02-13T12:15:59-0500

"mv=(m+m_1)u,"

"u=\frac{mv}{m+m_1}=3.1~\frac ms."